OpenStudy (anonymous):

Let u = <-8, -9>. Find 7u.

OpenStudy (nincompoop):

tell us what you know

OpenStudy (anonymous):

-8*7=56 -9*7=63 So it'd be <-56,-63>?

OpenStudy (anonymous):

Hello?